water is filled in a flask upto a height of 20 cm, the bottom of the flask has an area of 1m2 . If the atmospheric pressure is 1.01 × 105 Pa then what force is exerted by the water on the bottom? g= 10 m/s2 density of water = 1000 kg /m3 

Grade:10

3 Answers

Arun
5 years ago
Dear student
 
P = Po + rho* g* h
 
P = 1.01 * 10^5 + 0.20 * 1000*10
 
P = 1.01 *10^5 + 0.02 * 10^5
 
P = 1.03 * 10^5 N/m²
 
 
JC
5 years ago
Arun, I think u have given the wrong answer. The question asks for the force but u have given the pressure. To correct it, I suggest u to multiply pressure with the area to get the force. Hope it helps!
pavan
5 years ago
 
P = Po + rho* g* h
 
P = 1.01 * 10^5 + 0.20 * 1000*10
 
P = 1.01 *10^5 + 0.02 * 10^5
 
P = 1.03 * 10^5 N/m²
p=F/a
f=pxa
=1.03 * 10^5x3.14.10x10^-4
=3230N